### v1

# FragNet: A Graph Neural Network for Molecular Property Prediction with Four Layers of Interpretability

Abstract
Molecular property prediction is a crucial step in many modern-day scientific applications including drug discovery and energy storage material design. Despite the availability of numerous machine learning models for this task, we are lacking in models that provide both high accuracies and interpretability of the predictions. We introduce the FragNet architecture, a graph neural network not only capable of achieving prediction accuracies comparable to the current state-of-the-art models, but also able to provide insight on four levels of molecular substructures. This model enables understanding of which atoms, bonds, molecular fragments, and molecular fragment connections are critical in the prediction of a given molecular property. The ability to interpret the importance of connections between fragments is of particular interest for molecules which have substructures that are not connected with regular covalent bonds. The interpretable capabilities of FragNet are key to gaining scientific insights from the model’s learned patterns between molecular structure and molecular properties.
keywords:
Graph Neural Networks, Deep Learning, Attention, Model Interpretability
1 Introduction
Molecular property prediction is a crucial component of material discovery and design in many modern scientific applications including drug design, energy storage material discovery, catalysis and agrochemicals. Despite the availability of a large number of machine learning models[1] for molecular property prediction, there is often a trade-off in the ability of models to provide highly accurate predictions versus their ability to provide interpretability of their predictions to enable scientific insights. In this work, we introduce a graph neural network architecture called FragNet which is not only capable of achieving prediction accuracies comparable to or exceeding the accuracies of the current state-of-the-art models but also has the ability to provide insight into four levels of molecular substructures which are the atoms, bonds, fragments and fragment connections. In other words, with this model, one can understand which atoms, bonds, molecular fragments and also which connections between molecular fragments play critical role in predicting a given molecular property. By enhancing the model with the ability to reason about the connections between fragments, we provide improved representations for molecules with substructures that are not connected with regular covalent bonds, such as salts and complexes.
There are several prior works which implement deep learning models which leverage attention mechanisms which can provide values for different types of molecular substructures. The AttentiveFP [2] and MoGAT [3] models provide atom level importance values. The work by Wu et al. [4] provides fragment level attention. However, their work is limited in the range of molecular properties on which it evaluates, as it only provides results for the ESOL dataset [5] using a random train-test split configuration. In contrast, our work provides prediction results for several benchmark datasets in MoleculeNet [6] while using more challenging scaffold splitting method [7]. This provides a more comprehensive and robust understanding of how the reasoning of the molecular property prediction models is affected by the target property.
In addition to demonstrating the strong property prediction performance of the FragNet model across this challenging set of tasks, we also demonstrate the utility of the multi-layered interpretability mechanisms of the through several case studies on a selection of property prediction tasks. We use model attention weights and contribution values to investigate the reasoning used by the model for both individual molecular predictions as well as aggregated across multiple predictions, to identify the key molecular constituents that drive property variations. To further validate the reasoning extracted from the models, we perform a comparative study of FragNet contribution scores with density functional theory (DFT) computations of electrostatic surface potentials. Finally, we develop and release a interactive browser application to make these types of interpretability studies accessible for other molecular property tasks.
2 Results
2.1 Model
Figure 1: FragNet’s architecture and data representation. (a) Atom and Fragment graphs’ edge features are learned from Bond and Fragment connection graphs respectively. b) Initial fragment features for the fragment graph are the summation of the updated atom features that compose the fragment. (c) Illustration of FragNet’s message passing taking place between two non-covalently bonded substructures. Fragment-Fragment connections are also present between adjacent fragments in each non-covalently bonded structure of the compound.
Table 1: FragNet performance measured with RMSE on regression tasks in the MoleculeNet benchmark compared with four state-of-the-art baselines. Best results in bold. Lower is better. Reported here are mean $\pm$ standard deviation corresponding to three random seeds. Except for SimSGT, results for other models were obtained from the work of reference [8].
Dataset ESOL LIPO CEP
ContextPred[9, 8] 1.196 $\pm$ 0.037 0.702 $\pm$ 0.020 1.243 $\pm$ 0.025
AttrMask[9, 8] 1.112 $\pm$ 0.048 0.730 $\pm$ 0.004 1.256 $\pm$ 0.000
JOAO[10, 8] 1.120 $\pm$ 0.019 0.708 $\pm$ 0.007 1.293 $\pm$ 0.003
GraphMVP[11, 8] 1.064 $\pm$ 0.045 0.691 $\pm$ 0.013 1.2228 $\pm$ 0.001
Mole-BERT[8] 1.015 $\pm$ 0.030 0.676 $\pm$ 0.017 1.232 $\pm$ 0.009
SimSGT[12] 0.917 $\pm$ 0.028 0.670 $\pm$ 0.015 1.036 $\pm$ 0.022
FragNet 0.881 $\pm$ 0.011 0.682 $\pm$ 0.031 1.092 $\pm$ 0.031
Table 2: FragNet performance measured with mean $\pm$ standard deviation AUC-ROC corresponding to three random seeds on classification tasks in the MoleculeNet benchmark compared with four state-of-the-art baselines. Best result in bold. Higher is better.
Dataset Clintox Sider Tox21
ContextPred[8] 74.0 $\pm$ 3.4 59.7 $\pm$ 1.8 73.6 $\pm$ 0.3
JOAO[10, 8] 66.6 $\pm$ 3.1 60.4 $\pm$ 1.5 74.8 $\pm$ 0.6
AttrMask[9, 8] 73.5 $\pm$ 4.3 60.5 $\pm$ 0.9 75.1 $\pm$ 0.9
MGSSL[13, 8] 77.1 (4.5) 61.6 (1.0) 75.2 (0.6)
GraphMVP[11, 8] 79.1 $\pm$ 2.8 60.2 $\pm$ 1.1 74.9 $\pm$ 0.8
Mole-BERT[8] 78.9 $\pm$ 3.0 62.8 $\pm$ 1.1 76.8 $\pm$ 0.5
SimSGT[12] 85.7 $\pm$ 1.8 61.7 $\pm$ 0.8 76.8 $\pm$ 0.9
FragNet 86.8 $\pm$ 1.8 63.7 $\pm$ 1.9 76.9 $\pm$ 0.6
The FragNet model is based on a message-passing graph neural network architecture which reasons over four different graph-based representations of the molecular structure. An overview of the graph-based representations used by our model is shown in Figure 1. The data representation used by our model consists of four graph structures: atom-based, bond-based, fragment-based and fragment connection-based. The nodes and edges of the atom graph are atoms and covalent bonds respectively. In the bond graph, the nodes are bonds, while an edge in this graph is formed when two bonds have a common atom. The fragment graph is created by decomposing the molecules into substructures using a fragmentation scheme. In this work, we used BRICS fragmentation[14, 15, 16]. If the molecule is not a salt or a complex, the edges in the fragment graph are formed by bonds at which the fragments are split.
In the case of molecular complexes, where regular bonds do not exist between some fragments, we form ‘virtual’ connections to enable message passing between such disconnected substructures. Specifically, we create virtual connections between fragments that are linked by BRICS fragments. If a molecule cannot be decomposed into fragments, we treat the molecule itself as a fragment and establish a self-connection. For molecules composed of substructures that are not connected by covalent bonds, we create virtual connections by linking each fragment in one substructure to each fragment in the other.
Finally, the fragment-connection graph represents the connections between fragments as nodes with an edge formed when the connections have a fragment in common. These connections can be regular and/or virtual bonds depending on the type of the molecule. An important benefit of the fragment-connection graph is that it gives the ability to understand the interactions between different substructures of the molecules. This is particularly useful when regular bonds do not exist between two substructures.
The FragNet model leverages a hierarchical approach among these graph-representations, using the learned representations of lower-level structures to initialize the features of the high-level structures in the each subsequent graph representation. Our model starts by creating the representation for a given molecule using the bond graph. The nodes in the bond graph are initially featurized using bond properties, while the edges in the bond graph are featurized using bond angles. As bond properties we considered bond type (single, double, triple, aromatic), whether the bond is considered to be conjugated, whether the bond is in a ring and stereochemistry of the bond. The node representation of the bond graph is updated using the graph attention mechanism [17]. The updated node features of the bond graph are then used as the initial edge features of the atom graph, as indicated by the yellow arrow in Figure 1(a). Using these edge features, the atom graph is updated, also using the graph attention mechanism. Once updated, a molecular representation is created by summing all the atom feature vectors of the atom graph.
The fragment graph is updated following a similar procedure. For the initial node features of the fragment graph, we use the summed atom features from the atom graph corresponding to the atoms of which the fragment is composed. The fragment graph node representations are then updated through the graph attention mechanism. Then the edge features of the fragment graph are initialized using the learned node features of the fragment connection graph before applying graph attention. Using the updated fragment graph a second molecular representation is constructed. The final molecular representation is the concatenation of the representations constructed based on the atom and fragment graphs. This representation can then be used for downstream tasks like molecular property prediction. A complete mathematical description of the FragNet architecture is provided in the Methods section.
2.1.1 Prediction Accuracies
We test the performance of FragNet on multiple molecular property prediction tasks derived from the MoleculeNet benchmark including four regression tasks and three classification tasks. These tasks spanned a range of chemical, biological, and toxicity properties. Before performing property prediction training, the model was pre-trained on a set of self-supervised tasks as described in the Methods section. Table 1 and Table 2 presents the prediction accuracies of FragNet obtained based on scaffold splitting[18]. For each dataset, FragNet’s hyperparameters were optimized based on the validation accuracy using three random seeds. Using the optimized parameters and the same three random seeds, FragNet was fine-tuned, and the mean and standard deviation of the prediction accuracy on the test set are reported in Tables 1 and 2. We compare the performance of FragNet with four state-of-the-art baseline approaches. With the exception of CEP and Malaria, FragNet achieves accuracies comparable to or better than the current state-of-the-art for both regression and classification tasks. For CEP, the slightly lower accuracies may be attributed to fact that we only the first hyperparameter combination suggested by Optuna compared to thirty hyperparameter optimization runs conducted for the other datasets. This is due to the time constraints of training as the training set of CEP is large, with 23,982 molecules.
Figure 2: Different types of attention weights and contribution values available in FragNet visualized for CC[NH+](CCCl)CCOc1cccc2ccccc12.[Cl-] with atom, bond, and fragment attention weights shown in (a),(b), and (c) and fragment contribution values shown in (d). The top table provides the atom to fragment mapping and the bottom table provides the fragment connection attention weights. Atom and bond attention weights are scaled to values between 0 and 1. The fragment and fragment connection weights are not scaled. The numbers in blue boxes in (d) correspond to Fragment IDs in ‘Atoms in Fragments’ table.
2.2 Interpretability
The primary distinction between FragNet and other models lies in its ability to interpret predictions based on four different types of substructures. Notably, FragNet can handle molecules with substructures that are not connected via covalent bonds, such as salts and complexes. We can analyze all four substructure types using two different mechanisms from the model - attention weights and contribution values. Attention weights are derived from the graph attention mechanisms from each of the four graph representations and provide insight into which substructures the model focuses on when making its predictions. Meanwhile, we can quantify the contribution of a given substructure to the prediction of a molecular property value [4]. This is achieved by first predicting the property for the entire molecule in the usual manner. Subsequently, another prediction is made after masking the substructure of interest, meaning the node features of this substructure are excluded from the prediction process. If the prediction with the masked substructure is lower than that of the unmasked molecule, the masked substructure is considered property-increasing. Conversely, if the prediction with the masked substructure is higher, the masked substructure is deemed to reduce the property value. Formally, the contribution is defined as follows:
$Contribution=Property^{unmasked}-Property^{masked}$ (1)
2.2.1 Interpretation Case Study
We first demonstrate the interpretation capabilities of FragNet using a sample compound. In the subsequent sections, we analyse the insights that can be gained by using attention weights and contribution values using three prediction tasks: solubility, lipophilicity and cancer drug response. In the following discussions, we use the notation A(i) to denote an atom of type A with atom index i, and A(i)-B(j) to denote a bond connecting atoms A(i) and B(j).
The attention weights and contribution values of 1-Naphthyloxyethylethyl-beta-chloroethylamine hydrochloride from the solubility prediction model are illustrated in Figure 2. The largest atom attention weights are on the atoms in the vicinity of positively charged nitrogen, N(2) atom. This is expected as this region likely interacts most significantly with the solvent water molecules. The next largest weights are on four atoms in the aromatic ring and ether O(8). A significant attention on ether oxygen is also understandable as it can act as a hydrogen bond acceptor allowing it to interact with water molecules. These ether oxygen can also withdraw electrons from the adjacent -C2H2- group and donate them to double ring system[19].
In terms of bond attention weights, the largest weights are on the C(0)-C(1), C(3)-C(4), and C(6)-C(7) bonds. One atom of each of these bonds are connected to N(2)+. Note that these three bonds are part of fragments 0, 2, and 3. Examining the 3D structure of this molecule, we observe that these three fragments could cause steric hindrance to N(2)+. We hypothesize that the large weights on C(0)-C(1) and C(3)-C(4) could be attributed to two factors: (a) the steric effects caused by fragments 0, 2, and 3, and (b) the electron flow towards the N(2)+ atom, as alkyl groups generally exhibit electron-donating properties. It should also be noted that due to the presence of chlorine in fragment 2, its electron-donating capability might be less than that of fragment 0. Moreover, there could be an electron flow through C(6)-C(7) towards the double ring system because of the ether oxygen bonded to the fragment 3. Ethers are known to exhibit both electron-withdrawing and electron-donating effects[19]. In this molecule, it is possible that the ether oxygen is withdrawing electrons from fragment 3 and donating them to the double ring system. The large weight on the C(6)-C(7) bond could signify this interesting dynamic in electron flow, as it is situated between two electron-withdrawing atoms, O(8) and N(2).
The fragments containing N+ and Cl- are the ones with the highest attention weights. This result is not surprising, as these ionic species tend to interact strongly with solvent water molecules. The strongest connection weight is observed between the N+--H fragment and the --C2H2-- group adjacent to the ether oxygen. Although the exact reasons for this interaction remain unclear, it is possible that the --C2H2-- group contributes to the stability of the compound by shielding the N+ atom from the electron-withdrawing effects of the ether oxygen. Additionally, Cl- shows significant interaction with the double ring system, while the ether oxygen interacts robustly with its adjacent hydrocarbon group. Subsequent large interaction weights are noted between the ether oxygen and Cl-, N+ and its adjacent chloroethyl group, N+ and its adjacent ethyl group as well as between N+ and Cl-.
In Figure 2(d), we show the contribution values for the molecule fragments which provide information about the directional impact of different substructures on the predicted property. From the fragment contributions, we identify that the double ring system is the most hydrophobic based on its highly negative contribution value indicating that it significantly decreases the model’s predicted solubility. The next most hydrophobic groups are fragments 2 and 0. The hydrophobicity of these two fragments can be mainly attributed to the steric effects they impose on N+. Given their ability to form hydrogen bonds with solvent water, it is not surprising that the most hydrophilic substructures are Cl- and those containing oxygen and N+.
2.2.2 Substructure Analysis
In the following sections, we perform a quantitative study of the highly impactful substructures for predicting solubility, lipophilicity, and cancer drug response. To identify these substructures, we first filter the model predictions to those with low error to focus on cases where the model is correctly inferring the structure-property relationships. We then identify substructures which appear in multiple molecules and average their attention weights across all the predictions. Substructures which have high average attention across the predictions are likely to have a significant impact on the predicted property.
Figure 3: Mean atom attention weight values when compounds with prediction error $<$ 0.1 were considered. Error bars indicate the standard deviation.
Figure 4: Mean bond attention weight values when compounds with prediction error $<$ 0.1 were considered.Error bars indicate the standard deviation.
Figure 5: Fragment and fragment-fragment connection attention weights and contribution values corresponding to solubility prediction. Left: violin plots for attention weight distributions of highly weighted fragments and the corresponding contribution value distributions. Right: Fragments with largest and smallest contribution values corresponding to the fragments which tend to decrease and increase the prediction solubility respectively. The average attention weight value for each fragment is shown above each boxplot.
Figure 6: Fragment and fragment-fragment connection attention weights and contribution values corresponding to lipohilicity prediction. The contribution values corresponding to largest weights are also marked.
Atom and Bond Weights
Figure 3 illustrates the atoms with largest average attention weights for solubility, lipophilicity, and cancer drug response (CDR) prediction for molecules with absolute errors less than 0.1. It is observed that Nitrogen (N) and Carbon (C) play dominant roles in all three scenarios, followed by O and S. Compared to solubility and lipophilicity, F and Cl have a more significant impact on cancer drug response prediction.
Highly weighted bonds in compounds making absolute prediction errors less than 0.1 are shown in Figure 4. The significance of the carbon-carbon (C-C) bond is greater for solubility and lipophilicity prediction than for cancer drug response prediction. It is unsurprising that the carbon-oxygen (C-O) bond is more critical for solubility prediction than for lipophilicity prediction, primarily because of the polarity of the C-O bond. Polar substances are more soluble in polar solvents like water.
Fragment weights
Next, we study the attention weight and contribution score distributions for the fragment-level substructures. In the left panel of Figure 5, we present the attention weight distributions of highly weighted fragments in the context of solubility prediction. Meanwhile, the fragments associated with largest and smallest contribution values are shown in the right panel of Figure 5. These results are based on test set predictions with an absolute error below 0.1, and correspond to an $R^{2}$ score of 0.99 between observed and predicted solubility values.
We find that the fragments containing N are among the high weight fragments ($[*]N([*])[*]$) and high contribution value fragments ($CN(C)[*]$, $N[*]$). Nitrogen containing functional groups are known to be hydrophilic. Agreeing with the known chemistry, -OH which is also a well known hydrophilic functional group that has on average a positive contribution value.
The compounds containing $[Cl-]$ are salt structures. Our analysis reveals that Cl- is either the most significant or the second most significant fragment in all but one of the compounds containing Cl- in our study set according to the attention weights. Additionally, these compounds contain N+, which ranks among the two most significant substructures in 8 out of the 14 compounds. The fragment weights for these 14 compounds are illustrated in Figure S1. Considering the chemical properties, Cl- is likely to interact predominantly with the positively charged (N) atom. Consequently, the model’s identification of these structures as the most important underscores its focus on the relevant substructures.
Despite the presence of the polar C=O group, the substructures $O=C([*])[*]$ and $O=C([*])C[*]$ also have most of their contribution values between -0.5 and 0. This is likely because of the other fragments attached to these substructures. Upon further inspection, we found that the contribution of $O=C([*])[*]$ is positive when one of the substructures attached to it is OH or NH2 (see Figure S2). This result is consistent with the well-known fact that the carboxylic acid group (O=C(C)OH) is hydrophilic. When we inspected the compounds where $O=C([*])[*]$ has negative contributions, we find that in majority of them O is attached to one end of $O=C([*])[*]$ and a five membered ring system to the other end (structures are shown in Figure S3).
Figure 7: Fragments with (a) largest attention weights for cancer drug response prediction. (b) Fragments corresponding to five largest and five smallest contributions towards CDR prediction.
Figure 8: Attention weights (top panel) and contributions (bottom panel) for drugs with large and small drug response values.
Even though CN(C)[*] has the highest average contribution value, we see that the contributions of N[*] and O[*] can be even larger. As there are only a few such compounds in our study set, we currently cannot distinctly determine the differences between the compounds where $N[*]$ and $O[*]$ contribute greater than $CN(C)[*]$ for solubility. The substructures containing benzene-like rings are observed to be hydrophobic irrespective of the other structures they are attached to, as their contribution values are consistently negative. As a comparison to solubility, fragment attention weights and contribution values corresponding to lipophilicity are shown in Figure 6. As expected, the trends observed are opposite to those for solubility in terms of the types of substructures which enhance versus diminish the property.
The highly weighted fragments for cancer drug response prediction are shown in Figure 7(a). Among them are amino groups (–NH2, –NR2), carbonyl (C=O) groups, ether oxygens, hydroxyl groups and aromatic rings, which are known to be important building blocks of drug molecules [20, 19].
Because of their polarity, amino groups increase water solubility and participate in hydrogen bonding and ionic interactions with amino acid residues in proteins, nucleic acids, and other biomolecules. Studies have shown that a compound conjugating with an amino acid results in the improvement of its pharmacological activity[21, 22]) and water solubility[23, 24], and the reduction of its cytotoxicity[25].
The presence of a carbonyl (C=O) group can improve the molecular recognition process, allowing the drug to directly fit into the active sites of the target[26]. Carbonyl groups can enhance the stability of the drug molecule due its ability to form hydrogen bonds. This network of hydrogen bonds helps to prevent premature degradation and ensure that the drug reaches its site of action intact. The chemical reactivity of carbonyl groups can play a significant role in the mechanism of drug action as well[26].
Figure 9: Highly weighted connections in the contexts of solubility, lipophilicity and cancer drug response prediction.
Figure 10: 2D UMAP projection of FragNet embeddings for molecules in the PNNL solubility dataset highlighting that molecules with hydrophobic and hydrophilic groups occupy different regions of the latent space.
Apart from contributing to improve solubility, due to the electron donating and accepting ability of ether oxygens, the presence of ether groups can also increase the bioavailability of a drug[19]. Ether oxygens (-O-) can also bolster the stability of drug molecules, reducing the chances of degradation before the drug reaches its intended site of action. For instance, in penicillin V, the ether oxygen can attract the lone pair of electrons from the oxygen of the side chain carbonyl group. This withdrawal of electrons prevents the lone pair from attacking the carbonyl carbon of the $\beta$-lactam ring[19].
The presence of hydroyl (-OH) group can increase the solubility and thereby the oral absorption of drug molecules. Hydroxyl groups also affect the metabolism, duration and the selectivity[19]. Due to their ability to engage in hydrogen bonding, hydroxyl groups can be employed to modify the sites where hydrogen bonds are formed within a drug molecule. This capability allows for the regulation of both the binding affinity and specificity of the drug towards its biological target[20].
We also find fragments containing aromatic rings among the highly weighted. Aromatic rings provide hydrophobic interactions and stacking interactions with biological targets and are often present in drugs that need to pass through lipid membranes.
In Figure 7 (b), fragments associated with the five largest and smallest contribution values in our test dataset are shown. Notably, the hydrophobic fragments also contribute negatively to the cancer drug response prediction.
To underscore the significance of identifying useful substructures in cancer drug design, Figure 8 presents data on four drugs tested on the same BRCA[27, 28] cell line type. Among these, three drugs exhibit high response values (approximately 4.1 to 4.5), while the fourth drug shows a notably lower response value of around -0.66. Analysis reveals that the drugs with high response values have relatively smaller negative contribution values compared to the drug with the low response value. Specifically, the average contributions across fragments for the highly responsive drugs are -0.089, -0.094, and -0.095, while the low response drug has an average contribution of -0.24. Additionally, the highly responsive drugs contain fragments with substantially larger positive contributions compared to the low response drug. The positive contribution values for drugs (a), (b), and (c) are 0.40355, 0.16833, and (0.155, 0.020) respectively, whereas the low response drug shows positive contribution values of just 0.007, 0.007, and 0.001. This indicates that in computational drug design, FragNet’s contribution scoring can be utilized to identify initial fragments with positive contributions, thus aiding in the design of more effective cancer drugs.
Figure 11: Non Polar/Polar ratios (NPR) and contribution (C) values of fragments. It is observed that for highly hydrophobic fragments, the Non Polar Fraction takes large values compared to the hydrophilic fragments.
Figure 12: Interface of the interactive application
Fragment-Fragment Connection Weights
The highly weighted fragment-fragment connections for predicting solubility, lipophilicity, and drug response are shown in Figure 9. In most cases, these connections involve oxygen (O) or nitrogen (N) atoms. The presence of N-H connections is notably important for lipophilicity prediction. Additionally, C=O connections are also observed to be crucial for determining all three properties. In particular, the interaction between $[*]CC[*]$ and oxygen is significant for both solubility and drug response predictions. Non-covalent interactions between $CC[*]$ and substructures containing halogen atoms are also noteworthy, as observed in the solubility datasets. Upon close inspection, we found that in almost all cases, these ethyl groups are attached to electron-donating N or O atoms. Therefore, we believe these interactions signify non-covalent interactions between the electronegative group and the anionic substructure. Also, because of the electron donating nature of the ethyl groups, it is possible that there is an attraction between Cl- and slightly positively charged ethyl groups. This ethyl halogen interaction is also observed for when Cl- is the anion.
2.3 Additional Analysis
2.3.1 Model Embeddings
To further demonstrate the effectiveness of FragNet’s representations in distinguishing different classes of molecules, we present UMAP results in Figure 10 which show the 2D projection of the embeddings from the layer immediately preceding the final linear layer of the model. The embeddings from this layer should include a condensed representation of the molecular structure that encodes the information necessary for predicting the molecular property. Therefore, the groupings of molecules in this space provides insight into the reasoning used by the model to infer the target property. For solubility, we observe that some molecules containing hydrophilic and hydrophobic fragments are located in separate regions in the embedding space.
2.3.2 Validations using Density Functional Theory (DFT) calculations
To further investigate the interpretability of FragNet, we calculated the electrostatic surface potentials of some molecules that had the FragNet model trained on solubility predicted with high accuracy using Density Functional Theory (DFT). Regions with electrostatic surface potentials (absolute value) greater than 10 were designated as polar areas, while those with potentials less than 10 were classified as non-polar areas. For each molecule, the ratio of the non-polar area to the polar area (NPR) was then calculated for the fragments. A larger NPR indicates a fragment with lower expected polarity. Therefore, we expect fragments with high NPR values to have negative or small contribution values. We observe this trend is in some molecules, as shown in Figure 11. This observation is prominent in highly hydrophobic molecules. Specifically, in molecules where the total NPR is greater 1, the average fragment contribution is -1.19, and in molecules where the total NPR is less than 1 the average fragment contribution is -0.35. In Table S1, we provide the fragment contribution and NPR values for 16 molecules.
2.3.3 Interactive Browser Application
We are also releasing the code for a web browser-based application that permits users to edit molecules and examine how attention weights vary across different substructures (Figure 12). This tool is designed for educational purposes and small-scale molecule design. It supports fragmenting a molecule using both BRICS, Murcko, and BRICS+Murcko fragmentation schemes. Currently, users can visualize atom, bond, and fragment weights, as well as fragment contributions. Additionally, substructure weights and substructure-substructure connection weights are displayed.
3 Conclusion
We developed a new graph neural network (GNN) architecture capable of providing model interpretations using atoms, bonds, fragments, and fragment connections. To the best of our knowledge, this is the first GNN architecture that accounts for message passing between non-covalently bonded fragments, thereby creating an improved representation for salts and ionic liquids. The prediction accuracies of our model, FragNet, are comparable to the latest state-of-the-art models. Fragment masking allows for the determination of fragments that enhance or diminish property values, a capability particularly useful in molecular design applications. We find that the model correctly focuses on fragments that most affect solubility, lipophilicity, and drug response based on known chemical properties, thus validating the FragNet architecture. Our code is publicly available at github/fragnet, and we have also provided an interactive web-based application for editing molecules and visualizing the corresponding attention weights of the substructures.
4 Methods
FragNet Architecture
The importance of node j and the edge connecting nodes $i$ and $j$ to node $i$ is defined by $\alpha_{i,j}^{bond-graph}$ in Equation (2).
$\alpha_{i,j}^{bond-graph}=\frac{exp(LeakyRelU(M_{i,j}))}{\sum_{k\in N(i)}exp(LeakyRelU(M_{i,k}))},$ (2)
where,
$M_{i,j}=\mathbf{a}^{b}[\mathbf{W}^{b}h_{i}^{b}||\mathbf{W}^{b}h_{j}^{b}||\mathbf{W}^{be}e_{i,j}|.$ (3)
$h_{j}^{b}$ is a feature vector of a neighboring node of node $i$ in the bond graph. Note that these nodes correspond to bonds in the atom graph. $a^{b}$ is a trainable weight vector. $e_{i,j}$ is the feature vector for the edge connecting nodes $i$ and $j$ in the bond graph. These attention weights determine the level of impact that each other bond has $h_{j}^{b}$ in updating the learned representation of a given bond, $h_{i}^{b}$
The updated node features of the bond are calculated as,
$h_{i}^{bond-graph}=\sum_{j\in N(i)}\alpha_{i,j}^{bond-graph}\textbf{W}h_{j,0}^{bond-graph}$ (4)
We do not consider self edges in the bond graph, as it can significantly increase the computational cost. Hence $N(i)$ includes only the first nearest neighbors.
We then use these $h_{i}^{bond-graph}$ as the edge features of the atom graph. As in the case of bond graph, the attention coefficients of the atom graph will be computed considering both node features and edge features. The initial node features of the atom graph are based on the atomic properties like atom type, number of valence electrons, etc. For the message passing in the atom graph, we do consider self edges (the edges connecting an atom to itself). The edge properties of the self edges are initialized as a vector of zeros having the same dimension as $h_{i}^{bond-graph}$.
$h_{i}^{atom-graph}=\sum_{j\in N(i)}\alpha_{i,j}^{atom-graph}\textbf{W}h_{j,0}^{atom-graph}$ (5)
Our model also considers a graph of molecular fragments. We use BRICS fragmentation to identify the fragments. In the case of regular molecules (not salts or complexes), the edge between two molecular fragments is a regular covalent bond. In the case of molecules with isolated fragments (like salts and complexes), where regular covalent bonds do not exist between some fragments, we create virtual bonds connecting each such isolated fragment to every other fragment (illustrated in Figure 1 [c] ). There are some molecules which cannot be fragmented. For such molecules we create a self-connection connecting itself to itself to facilitate message passing.
The initial fragment features are calculated by summing the features of the atoms belonging to the fragments.
$h_{i,0}^{fragment-graph}=\sum_{j\in Atoms\_of\_frag_{i}}h_{j}^{atom-graph}$ (6)
We then use graph attention mechanism to update fragment features $h_{i,0}^{fragment-graph}$ similar to how we updated atom features of the atom graph. In order to achieve this task, we also have to consider a graph of connections, similar to the bond graph. The node features of the connection graph are one-hot encoded connection types. This connection type could be single, double, triple, aromatic, ‘self connection’, or ‘virtual connection’. The edge attributes of the connection graph are obtained by summing the node features of the nodes at the either side of the edge. The graph attention coefficients for the connection graph are computed using a expression similar to Equation 2. Then these coefficients are used to update the node features of the connection graph, which are the connections of the fragment graph. The updated node features of the fragment connection graph are used to initialize the edge features of the fragment graph.
Features
The atomic and bond features were created using the methods implemented in the RDKit package[29].
As atomic features we consider Atomic Number, Implicit Valence, Formal Charge, Number of Radical Electrons’, Hybridization, Is the atom Aromatic, Is the atom in a Ring, Total Number of H atoms, Chirality.
As the bond features we consider the Bond Type (SINGLE, DOUBLE, TRIPLE, AROMATIC), Is the bond Conjugated, Is the bond in a Ring, Stereo nature (STEREOANY, STEREOZ, STEREOE, STEREONONE), and Bond Direction (BEGINWEDGE, BEGINDASH, ENDDOWNRIGHT,ENDUPRIGHT, NONE).
Data
We leverage several datasets from MoleculeNet[18] to evaluate the performance of the FragNet model. The datasets include ESOL[5], Lipophilicity, Clintox[30, 31, 32], SIDER[33, 34] and Tox21. CEP dataset was obtained following the instructions given by [35]
For substructure studies, we used the PNNL solubliity dataset, the Lipophilicity dataset, and the CDRP dataset. The Lipophilicity dataset is the same that was used for model performance evaluation downloaded from MoleculeNet. The PNNL solubility dataset is a combination of [36], Cui[37] and SOMAS[38] datasets. The CDRP dataset was obtained from the work of [39].
Pretraining and Finetuning
We pretrained FragNet using 2,054,100 molecules from the dataset used by Uni-Mol[40]. The pretraining is based on the method outlined in the reference [41] and focuses on minimizing losses based on the predictions of bond length, bond angle, dihedral angle, and conformational energy of the molecule.
Hyperparameter tuning was performed using the Optuna package [42]. For small datasets we used a regression head with four layers. The number of nuerons in each layer were allowed to take values from 64 to 2048 in steps of 64. We also tuned the activation function (with options relu, silu, gelu, celu, selu, rrelu, relu6, prelu, leakyrelu) and batch size (with options 16, 32, 64, 128).
DFT calculation
Density functional theory (DFT) calculations were performed using the B3LYP[43, 44] functional with ORCA package.[45] The geometries of the molecules were optimized with def2-SVP basis set. Single point energies were calculated with def2-TZVP basis set. Vibrational frequencies were calculated for validation of stable configuration. An effect of implicit solvent model with water was included via Conductor-like Polarizable Continuum Model (CPCM) model.[46] The surface charge potentials were calculated by Multiwfn software.[47]
Declarations
Funding
This work was supported by Energy Storage Materials Initiative (ESMI), which is a Laboratory Directed Research and Development Project at Pacific Northwest National Laboratory (PNNL). PNNL is a multiprogram national laboratory operated for the U.S. Department of Energy (DOE) by Battelle Memorial Institute under Contract no. DE- AC05-76RL01830.
Conflict of interest
There are no conflicts of interest to declare
Consent for publication
Not applicable
Data availability
Except for the PNNL solubility dataset, the other datasets are publicly available. The PNNL solubility dataset was created by combining three publicly available datasets. The script to create the PNNL dataset is provided in the code’s github repository.
Code availability
The code is available at https://github.com/pnnl/FragNet
Author contribution